Chapter 1

An introduction to strongly correlated electron systems. Why there are interesting. The role play by the Mott transition

.

Renormalization group perspective. Effective hamiltonians

Chapter 3

Effective Action Approach Probes of the solid state. Photoemission Spectra. Other probes of the solid state

Chapter 5

Density Functional Theory. Kohn Sham Representation.

LMTO’s. ASA and full potential. Application: tight binding

.

Chapter 7

Hartree Fock

.

Chapter 8 LDA+ U

15

Chapter 9

Slave Bosons and Gutzwiller approximation.

Chapter 10

Mean Field Theory Classical Systems

Chapter 11

Mean Field Theory Quantum Systems

Chapter 12

Lattice models and their zero dimensional counterparts.

Chapter 13

Some simple examples. The falikov Kimball model

Chapter 14

Introduction to the Anderson IMpurity model.

Chapter 15

Anderson Impurity model. Equations of motion.

Anderson Impurity model. Second Order Perturbation Theory.

Chapter 17

Anderson Impurity model. Exact Diagonalization.

Chapter 18

Anderson Impurity model. Quantum Montecarlo.

Chapter 19

Anderson Impurity model. Orbitally Degenerate Case

.

Chapter 20

Mott transition. General arguments for its existence.

Chapter 21

Mott transition: Hubbard bands.

Chapter 22

The Friedel -Luttinger Theorem.

Results from the Mott transition. Half filled. Transfer of spectral weithg.

Chapter 24

LDA+ DMFT.

Chapter 25

Some interesting strongly correlated materials.

Chapter 26

Transition metals Iron and Nickel.

Chapter 27 Actinides.

53

Chapter 28

LaSrTiO3 3-d doped Mott insulator.

Chapter 29

Copper Oxides